OpenStudy (anonymous):

Given u--> = (1,-1,3) and v--> = (-4,5,3) find an ordered triple that represents 3u--> - 2v-->

OpenStudy (whpalmer4):

This is just like the distributive property. 3u = 3(1,-1,3) = (3, -3, 9) -2v = -2(-4,5,3) = ? 3u + -2v = (add the corresponding components)

OpenStudy (anonymous):

ohhhh I get it cx